ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

16. Applications

1991 MS Classification: 92 Yamashita, Hajime, Waseda University, Japan Approximate Analysis of Fuzzy Graph and its Application It is important to clarify the global feature of a fuzzy graph F if we analyse the inexact information such as the mental process and cognition. For this purpose, we define the p-graph F^{p} = (f_{ij}^{p}), f_{ij}^{p} = 1(f_{ij} > 0.5 = p), f_{ij}^{p} = 0(f_{ij} < 0.5 − p), f_{ij}^{p} = f_{ij} (otherwise), and the relative Humming distance d(F, F^{p}) = 2 ∑ | f_{ij} − f_{ij}^{p} | /(n^{2} − n). If p ↑ 0.5, then d(F, F^{p}) ↓ 0, so we have an approximate graph F^{∗} = F^{p} when it satisfies d(F, F^{p}) < ε. This approximate graph is used for the cognition analysis in education, the socimetry analysis in psychology and so on. Reference: H. Yamashita, et al, Approximate Expression of a Fuzzy Graph and its Application, 7th IFSA Congress (1997), 435 - 438.